How many atoms of nitrogen are in 10.0g of N2O5 ?
1
Expert's answer
2020-05-01T14:40:52-0400

The molecular weight of N2O5 equals:

Mr(N2O5) = (2 × 13 g/mol) + (5 × 16 g/mol) = 108 g/mol

The number of moles of 10 g of N2O5 euals:

n = m / Mr = 10 g / 108 g/mol = 0.09 mol

As 1 mole of any substance contains 6.022 × 1023 particles, 0.09 mol of N2O5 contains:

N(N2O5) = 0.09 mol × 6.022 × 1023 = 5.4 × 1022 molecules.

As one molecule of N2O5 contains two atoms of nitrogen:

N(N) = 5.4 × 1022 × 2 = 1.1 × 1023 atoms of nitrogen


Answer: 1.1 × 1023 atoms of nitrogen